At least one of these answers are wrong (probably more). Please help me get the right answer with the correct rounding! Thank you.McGilla Golf has decided to sell a new line of golf clubs. The clubs will sell for $840 per set and have a variable cost of $
Answer #1

1.
=(28980000+1340000)/((58000*(840-440)-9900*(1140-740)+11400*(480-250)-9140000-28980000/7)*(1-38%)+28980000/7)=3.20478942673166

2.
=-28980000-1340000+1340000/1.1^7+((58000*(840-440)-9900*(1140-740)+11400*(480-250)-9140000-28980000/7)*(1-38%)+28980000/7)/10%*(1-1/1.1^7)=16426963.3656112

3.
=RATE(7,((58000*(840-440)-9900*(1140-740)+11400*(480-250)-9140000-28980000/7)*(1-38%)+28980000/7),-28980000-1340000,1340000)=24.8287247775495%